# On quasiconformal non-equivalence of gasket Julia sets and limit sets

### v1

@@ 1. Introduction
Figure 1.1. The Apollonian gasket on the left, and the Julia set that is homeomorphic to an Apollonian gasket on the right. The two sets are not quasiconformally homeomorphic as Fatou components touch at an angle on the right.
@@ 1.1. Historical Background
  • the quasisymmetry group $\mathcal{QS}(\mathcal{J})$ equals to the homeomorphism group $\operatorname{Homeo}(\mathcal{J})$.
@@ 1.2. Strategy and techniques
We also recall that a quadratic polynomial $f(z)=z^{2}+c$ is called a Misiurewicz polynomial if the critical point 00 is strictly pre-periodic. The Julia set of a Misiurewicz polynomial is a dendrite. We remark that not every Misiurewicz polynomial $f$ is matable with the fat basilica. In fact, $f$ is matable with the fat basilica if and only if $f$ is not in the 1/2-limb of the Mandelbrot set [Tan92]. We also remark even if $f$ is matable with the fat basilica, the Fatou graph of their mating may not be a tree. In this case, by Theorem 1.6, the mating is not a fat gasket Julia set.
@@ 1.4. Structure of the paper
We study the Kleinian groups with gasket limit set in §2. The induced dynamics on the Fatou graph of a rational map with a fat gasket Julia set is studied in §3, where Theorem 1.5 is proved. The realization of rational maps with fat gasket Julia sets are studied in §4. In particular, Theorem 1.6 is proved there. Finally, the combinatorics of Fatou graphs for quadratic fat gasket Julia sets are studied in §5, and Theorem 1.3 is proved in §6.
@@ 2. Kleinian groups with gasket limit set
In this section, we recall some results on geometrically finite Kleinian groups, especially those with gasket limit sets. We refer to [LZ23, §8] for details. Throughout the section, let $\Gamma\subset\operatorname{PSL}(2,\mathbb{C})$ be a Kleinian group, $\Lambda\subset\hat{\mathbb{C}}$ its limit set, and $\Omega=\hat{\mathbb{C}}-\Lambda$ its domain of discontinuity. Up to a finite-index subgroup, we may assume $\Gamma$ is torsion-free. We also denote by $M=\Gamma\backslash\mathbb{H}^{3}$ the corresponding hyperbolic three-manifold.
@@ 2.1. Geometric finiteness and acylindricity
$\operatorname{core}(M):=\Gamma\backslash\operatorname{Cvx\,Hull}(\Lambda),$
@@
One can recognize acylindricity from the limit set in the geometrically finite case. The following characterization is well-known; see for example [LZ23, Prop. 8.4].
@@
One gradient in the proof is the following lemma, essentially from [Mas74, Theorem 3].
@@
    Let $\sigma_{i}$ be a curve on $X_{i}$ that is not null-homotopic in $\overline{M}:=\Gamma\backslash\{\mathbb{H}^{3}\cup\Omega\}$. Then $\sigma_{1}$ and $\sigma_{2}$ are not homotopic in $\overline{M}$.
@@ 3. Induced dynamics on the Fatou graph
Since vertices in $\mathcal{G}$ are in bijective correspondence with Fatou components of $f$, $f$ naturally induces a map $f_{*}$ on the vertex set of $\mathcal{G}$. By Lemma 3.2, there are no critical points on the boundary of a Fatou component. Thus, if $v,w$ are adjacent in $\mathcal{G}$, then $f_{*}(v)\neq f_{*}(w)$. Hence, $f_{*}(v)$ and $f_{*}(w)$ are adjacent. Therefore, we can extend the map $f_{*}$ to $\mathcal{G}$ by sending an edge $E=[v,w]$ to the edge $[f_{*}(v),f_{*}(w)]$. By construction, the induced map is simplicial.
@@ 4.2. Thurston theory of rational maps
where $\sigma\in\Sigma$, and $[\sigma^{\prime}]_{\Sigma}$ denotes the element of $\Sigma$ isotopic to $\sigma^{\prime}$, if it exists. The curve system is called irrreducible if $f_{\Sigma}$ is irreducible as a linear transformation. It is said to be a Thurston obstruction if the spectral radius $\lambda(f_{\Sigma})\geq 1$.
@@
Let $F:\mathcal{G}\longrightarrow\mathcal{G}$ be a degree $d$ simplicial branched covering satisfying the conditions in Theorem 4.1. Let $\widetilde{F}$ be the corresponding Thurston map on $-^\mathbb{S}^{2}\cong\hat{\mathbb{C}}$. Then $\widetilde{F}$ is equivalent to a rational map $f$.
@@
To get a fat gasket from the post-critically finite rational map $f$, we can perform a standard simple pinching deformation (see [CT18] or [HT04]) to create a double parabolic fixed point at the common boundary of the two fixed Fatou components of $f$. Let us denote this new map $\tilde{f}$. Note that $\tilde{f}$ is geometrically finite, and $\tilde{f}:J(\tilde{f})\longrightarrow J(\tilde{f})$ is topologically conjugate to $f:J(f)\longrightarrow J(f)$.
By the local Fatou coordinate at the double parabolic fixed point, it is easy to see the two fixed Fatou of $\tilde{f}$ are tangent to each other. By inductively pulling back this tangent point, we see that $J(\tilde{f})$ is a fat gasket. Since $\tilde{f}$ and $f$ have the same Julia dynamics, the induced dynamics of $\tilde{f}$ on the Fatou graph is conjugate to $F:\mathcal{G}\longrightarrow\mathcal{G}$. ∎
@@ 5.1. The canonical finite core and critical loop
Let $\mathcal{C}$ be the critical loop. Then $\mathcal{C}$ contains $E_{0}$ and the critical vertices of $\mathcal{G}$. Moreover, $f(\mathcal{C})$ is simple path of length $l$, and $f$ maps each of the two components of ^$\hat{\mathbb{C}}-\mathcal{C}$ homeomorphically to ^$\hat{\mathbb{C}}-f(\mathcal{C})$.
@@ 5.2. Simple closed curves in $\mathcal{G}$
Otherwise, $\gamma$ and $\mathcal{C}$ bounds some simple closed curve $\mathcal{K}$. Since $\mathcal{C}=f^{-1}(f(\mathcal{C}))$, $f(\mathcal{K})$ contains a simple closed curve. Thus, inductively applying Lemma 5.4, we can conclude that there exists some vertex $v$ of $\mathcal{K}$ and some iterate $j\geq 1$ so that $f^{j}(v)=c$. Since no vertex in $\mathcal{C}$ is mapped to $c$ by $f^{j}$ for any $j\geq 1$, together with the previous claim, there exists $j$ so that $f^{j}(v_{k-1})=c$. Therefore, $d(f(a),c)=k-1$.
Let $\alpha^{\prime}$ be a path of length $k-1$ connecting $f(a)$ and $c$, and let $\alpha=\alpha^{\prime}\cup E_{0}$. Note that $\alpha$ is a path of length $k$ connecting the two critical points that contains $E_{0}$. Denote the vertices of $\alpha$ in order by $w_{0}=a,w_{1},...,w_{k}=c$. Since $f(w_{1})=a$, we have that for any $i\leq k-1$ and any $j\geq 1$
@@ 5.3. Shortest anchored simple closed curves
  • if $f$ is Type IIB, then the critical loop has two siblings, and other than these three, each has exactly one sibling.
@@
Case 1: If $f$ is Type I, then $\mathcal{C}\cap f(\mathcal{C})=E_{0}$. Note that $f$ maps each of the two components of ^$\hat{\mathbb{C}}-\mathcal{C}$ homeomorphically to ^$\hat{\mathbb{C}}-f(\mathcal{C})$. Thus, by pulling back, it is easy to verify that $f^{-1}(\mathcal{C})$ is a figure eight curve, and one component, denoted by $\mathcal{C}_{1}$, is a shortest anchored simple closed curve. Since $\mathcal{C}\cap f(\mathcal{C})=E_{0}$, we have $\mathcal{C}\cap\mathcal{C}_{1}=E_{0}$, so they are not siblings. Similarly, $f^{-1}(\mathcal{C}_{1})$ is a figure eight curve, and contains a shortest anchored simple closed curve, denoted by $\mathcal{C}_{1}$. Note that $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$ are in two sides of the critical loop $\mathcal{C}$ (see Figure 5.2).
@@
Case 2: If $f$ is Type IIA, then $E_{0}\subsetneq\mathcal{C}\cap f(\mathcal{C})\subsetneq f(\mathcal{C})$. Similar as in Case 1, $f^{-1}(\mathcal{C})$ is a figure eight curve, and contains a shortest anchored simple closed curve, denoted by $\mathcal{C}_{1}$. Since $\mathcal{C}\cap f(\mathcal{C})$ contains more than $E_{0}$, $\mathcal{C}$ and $\mathcal{C}_{1}$ are siblings (see Figure 5.3). Similarly, $f^{-1}(\mathcal{C}_{1})$ contains one shortest anchored simple closed curve, denoted by $\mathcal{C}_{2}$. Note that $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$ are in two sides of the critical loop $\mathcal{C}$ and $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$ are not siblings. Then the same pull back argument as in Case 1 gives a sequence of shortest anchored simple closed curves $\mathcal{C}_{n}$, where $f(\mathcal{C}_{n})=\mathcal{C}_{n-1}$, and $\mathcal{C}_{n}$ has no sibling for all $n\geq 2$.
Case 3: If $f$ is Type IIB, then $\mathcal{C}\cap f(\mathcal{C})=f(\mathcal{C})$. Note that in this case, $\mathcal{C}$ contains both critical values. Thus, $f^{-1}(\mathcal{C})$ is a union of four simple paths connecting the critical points. Note that $f$ maps each one of the four components of ^$\hat{\mathbb{C}}-f^{-1}(\mathcal{C})$ homeomorphically to a component of ^$\hat{\mathbb{C}}-\mathcal{C}$. It is easy to verify that $f^{-1}(\mathcal{C})$ contains three shortest anchored simple closed curves, one of them is the critical loop. Denote them by $\mathcal{C}_{1},\mathcal{C}_{-1}$ and $\mathcal{C}$. Therefore, $\mathcal{C}$ has two siblings (see Figure 5.4).
Let $U_{1}$ and $U_{-1}$ be the regions bounded by $\mathcal{C}_{1}$ and $\mathcal{C}_{-1}$ that do not contain the critical loop $\mathcal{C}$, and let $\mathcal{V}_{1}$ and $\mathcal{V}_{-1}$ be the regions bounded by $\mathcal{C}$ that contains $\mathcal{C}_{1}$ and $\mathcal{C}_{-1}$ respectively. Then $f:U_{1}\longrightarrow V_{-1}$ and $f:U_{-1}\longrightarrow V_{1}$ are homeomorphisms. Thus, by inductively pulling back, we get a sequence of shortest anchored simple closed curves $\mathcal{C}_{n},n\in\mathbb{Z}$, where $f(\mathcal{C}_{n})=\mathcal{C}_{-\operatorname{sgn}(n)(|n|-1)}$ for $|n|\geq 1$. Here $\operatorname{sgn}(n)$ represents the sign of $n$, and $\mathcal{C}_{0}=\mathcal{C}$. Moreover, $\mathcal{C}_{n}$ has no sibling for all $|n|\geq 2$. ∎
@@ 6. Quasiconformal non-equivalence between quadratic gasket Julia set and limit set
A path $\gamma=[x_{0},...,x_{k}]$ in $\mathcal{G}$ is called a local geodesic if $x_{i},x_{j}$ are not adjacent in $\mathcal{G}$ if $|i-j|\geq 2$. Note that in particular, if $x,y$ are adjacent in $\mathcal{G}$, then there are no other local geodesics other than the edge $[x,y]$. We also remark that given any path $\gamma$, one can construct a local geodesic by replacing $[x_{i},x_{i+1},...,x_{j}]$ with $[x_{i},x_{j}]$ if $x_{i},x_{j}$ are adjacent in $\mathcal{G}$.
@@
  • $\gamma$ connects $a_{l}$ to $b_{i}$ for some $i=1,...,2l-2$;
@@
We claim that $\operatorname{Int}\gamma\cap f(\mathcal{C}_{0})=\emptyset$. Suppose not. Then we there exists a strictly shorter arc $\delta$ connecting $a_{l}$ to $v\in f(\mathcal{C}_{0})$. Note that $v\notin\{a,b\}$. Since $f$ maps the gap $R_{0}$ homeomorphically to its image, we can find a lift $\delta^{\prime}$ that connects $b_{l}$ and $a_{j}$ for some $j=1,...,2l-2$. Then $\tau(\delta^{\prime})$ is a local geodesic that connects $a_{l}$ to $b_{j}$ with $\operatorname{Int}\tau(\delta^{\prime})\subseteq\operatorname{Int}R_{0}$ and no interior vertex of $\tau(\delta^{\prime})$ is adjacent to $a_{0}$. This is a contradiction to the minimality of $\gamma$.
@@
separates the critical values $f(a_{l})$ and $b_{0}=f(a_{0})$ (see Figure 6.1). Since the arc $[a_{0},a_{1}]\cup...\cup[a_{l-1},a_{l}]$ has a lift connecting $b_{l}$ and $b_{0}$, we conclude that the arc $\alpha$ has a lift connecting $b_{l}$ and $a_{1}$. ∎
@@
Suppose that $f(\beta)\nsubseteq\overline{R_{0}}$. Since $\beta$ connects $b_{l}$ to $b_{0}$ in $R_{0}$, and $f$ sends locally the region bounded between $[b_{1},b_{0}]$ and $[b_{0},a_{2l-2}]$ in $R_{0}$ to the region bounded by $[a_{1},a_{0}]$ and $[a_{0},f(a_{2l-2})]$ in $R_{0}$ (see Figure 6.1), we can decompose $f(\beta)$ as $f(\beta)=\gamma_{1}\cup\gamma_{2}\cup\gamma_{3}$ where $\gamma_{1}$ connects $a_{l}$ to $b_{i}$ for some $i$, $\gamma_{2}$ connects $b_{i}$ to $b_{j}$ for some $j$ and $\gamma_{3}$ connects $b_{j}$ to $a_{0}$. Moreover, $\operatorname{Int}\gamma_{1},\operatorname{Int}\gamma_{3}\subseteq\operatorname{Int}R_{0}$.
@@
Otherwise, let $\beta_{1}\subseteq\beta$ be the preimage of $\gamma_{1}$ in $\beta$. Since $\beta$ is a local geodesic connecting $b_{l}$ to $b_{0}$, no interior vertex of $\beta_{1}$ is adjacent to $b_{0}$. Since $f^{-1}(a_{0})=\{b_{0},a_{1}\}$, there exists some interior vertex $x\in\beta_{1}$ that is adjacent to $a_{1}$. Consider the truncation $\beta_{1}^{\prime}\subseteq\beta_{1}$ that connects $b_{l}$ and $x$, and $\widetilde{\beta_{1}}=\beta_{1}^{\prime}\cup[x,a_{1}]$. Then $\tau(\widetilde{\beta_{1}})$ satisfies that
@@
  • $\tau(\widetilde{\beta_{1}})$ connects $a_{l}$ to $b_{1}$;
  • no interior vertex of $\tau(\widetilde{\beta_{1}})$ is adjacent to $a_{0}$.
@@
By Proposition 5.3, any simple closed curve in $\mathcal{G}$ has length $\geq 2l$. Thus, for any two vertices $v,w$, there exists at most one path connecting $v,w$ with length $<l$. Moreover, if there is a path connecting $v,w$ with length $l$, then all the other paths have length $\geq l$. Since $b_{i},b_{j},a_{0}$ all lie on a shortest loop $\mathcal{C}_{1}$, we have that the lengths $l(\gamma_{2}),l(\gamma_{3})\geq l$. So

### v2

@@ 1. Introduction
Figure $1.$ The Apollonian gasket on the left, and the Julia set that is homeomorphic to an Apollonian gasket on the right. The two sets are not quasiconformally homeomorphic as Fatou components touch at an angle on the right.
@@ 1.1. Historical Background
  • the quasisymmetry group $\mathcal{QS}(\mathcal{J})$ equals the homeomorphism group $\operatorname{Homeo}(\mathcal{J})$.
@@ 1.2. Strategy and techniques
We also recall that a quadratic polynomial $f(z)=z^{2}+c$ is called a Misiurewicz polynomial if the critical point 0 is strictly pre-periodic. The Julia set of a Misiurewicz polynomial is a dendrite. We remark that not every Misiurewicz polynomial $f$ is matable with the fat basilica. In fact, $f$ is matable with the fat basilica if and only if $f$ is not in the 1/2-limb of the Mandelbrot set [Tan92]. We also remark even if $f$ is matable with the fat basilica, the Fatou graph of their mating may not be a tree. In this case, by Theorem 1.6, the mating is not a fat gasket Julia set.
@@ 1.4. Structure of the paper
We study Kleinian groups with gasket limit sets in §2. The induced dynamics on the Fatou graph of a rational map with a fat gasket Julia set is studied in §3, where Theorem 1.5 is proved. The realization of rational maps with fat gasket Julia sets are studied in §4. In particular, Theorem 1.6 is proved there. Finally, the combinatorics of Fatou graphs for quadratic fat gasket Julia sets are studied in §5, and Theorem 1.3 is proved in §6.
@@ 2. Kleinian groups with gasket limit set
In this section, we recall some results on geometrically finite Kleinian groups, especially those with gasket limit sets. We refer to [LZ23, Appx. B] for details. Throughout the section, let $\Gamma\subset\operatorname{PSL}(2,\mathbb{C})$ be a Kleinian group, $\Lambda\subset\hat{\mathbb{C}}$ its limit set, and $\Omega=\hat{\mathbb{C}}-\Lambda$ its domain of discontinuity. Up to a finite-index subgroup, we may assume $\Gamma$ is torsion-free. We also denote by $M=\Gamma\backslash\mathbb{H}^{3}$ the corresponding hyperbolic three-manifold.
@@ 2.1. Geometric finiteness and acylindricity
[…]
@@
One can recognize acylindricity from the limit set in the geometrically finite case. The following characterization is well-known; see for example [LZ23, Prop. B.2].
@@
One ingredient in the proof is the following lemma, essentially from [Mas74, Theorem 3].
@@
[…]
@@ 3. Induced dynamics on the Fatou graph
Since vertices in $\mathcal{G}$ are in bijective correspondence with Fatou components of $f$, $f$ naturally induces a map $f_{*}$ on the vertex set of $\mathcal{G}$. By Lemma 3.2, there are no critical points on the boundary of a Fatou component. Thus, if $v,w$ are adjacent in $\mathcal{G}$, then $f_{*}(v)\neq f_{*}(w)$. Hence, $f_{*}(v)$ and $f_{*}(w)$ are adjacent. Therefore, we can extend the map $f_{*}$ to $\mathcal{G}$ by sending an edge $E=[v,w]$ to the edge $f_{*}(v),f_{*}(w)]$. By construction, the induced map is simplicial.
@@ 4.2. Thurston theory of rational maps
where $\sigma\in\Sigma$, and $\sigma^{\prime}]_{\Sigma}$ denotes the element of $\Sigma$ isotopic to $\sigma^{\prime}$, if it exists. The curve system is called irrreducible if $f_{\Sigma}$ is irreducible as a linear transformation. It is said to be a Thurston obstruction if the spectral radius $\lambda(f_{\Sigma})\geq 1$.
@@
Let $F:\mathcal{G}\longrightarrow\mathcal{G}$ be a degree $d$ simplicial branched covering satisfying the conditions in Theorem 4.1. Let $\widetilde{F}$ be the corresponding Thurston map on $\mathbb{S}^{2}\cong\hat{\mathbb{C}}$. Then $\widetilde{F}$ is equivalent to a rational map $f$.
@@
[…]
By the local Fatou coordinate at the double parabolic fixed point, it is easy to see the two fixed Fatou of $\tilde{f}$ are tangent to each other. By inductively pulling back this tangent point, we see that J(f~)J($\tilde${f}) is a fat gasket. Since $\tilde{f}$ and $f$ have the same Julia dynamics, the induced dynamics of $\tilde{f}$ on the Fatou graph is conjugate to $F:\mathcal{G}\longrightarrow\mathcal{G}$. ∎
@@ 5.1. The canonical finite core and critical loop
Let $\mathcal{C}$ be the critical loop. Then $\mathcal{C}$ contains $E_{0}$ and the critical vertices of $\mathcal{G}$. Moreover, $f(\mathcal{C})$ is simple path of length $l$, and $f$ maps each of the two components of $\hat{\mathbb{C}}-\mathcal{C}$ homeomorphically to $\hat{\mathbb{C}}-f(\mathcal{C})$.
@@ 5.2. Simple closed curves in $\mathcal{G}$
Otherwise, $\gamma$ and $\mathcal{C}$ bounds some simple closed curve $\mathcal{K}$. Since $\mathcal{C}=f^{-1}(f(\mathcal{C}))$, $f(\mathcal{K})$ contains a simple closed curve. Thus, inductively applying Lemma 5.4, we can conclude that there exists some vertex $v$ of $\mathcal{K}$ and some iterate $j\geq 1$ so that $f^{j}(v)=c$. Since no vertex in $\mathcal{C}$ is mapped to $c$ by $f^{j}$ for any $j\geq 1$, together with the previous claim, there exists $j$ so that $f^{j}(v_{k-1})=c$. Therefore, d(f(a),c)=k1d(f(a),c)=k-1.
Let $\alpha^{\prime}$ be a path of length k1k-1 connecting $f(a)$ and $c$, and let $\alpha=\alpha^{\prime}\cup E_{0}$. Note that $\alpha$ is a path of length $k$ connecting the two critical points that contains $E_{0}$. Denote the vertices of $\alpha$ in order by $w_{0}=a,w_{1},...,w_{k}=c$. Since $f(w_{1})=a$, we have that for any $i\leq k-1$ and any $j\geq 1$
@@ 5.3. Shortest anchored simple closed curves
  • if $f$ is Type IIB, then the critical loop has two siblings, each of which has exactly three siblings, and every other shortest anchored simple closed curve has exactly two siblings.
@@
Case 1: If $f$ is Type I, then $\mathcal{C}\cap f(\mathcal{C})=E_{0}$. Note that $f$ maps each of the two components of $\hat{\mathbb{C}}-\mathcal{C}$ homeomorphically to $\hat{\mathbb{C}}-f(\mathcal{C})$. Thus, by pulling back, it is easy to verify that $f^{-1}(\mathcal{C})$ is a figure eight curve, and one component, denoted by $\mathcal{C}_{1}$, is a shortest anchored simple closed curve. Since $\mathcal{C}\cap f(\mathcal{C})=E_{0}$, we have $\mathcal{C}\cap\mathcal{C}_{1}=E_{0}$, so they are not siblings. Similarly, $f^{-1}(\mathcal{C}_{1})$ is a figure eight curve, and contains a shortest anchored simple closed curve, denoted by $\mathcal{C}_{1}$. Note that $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$ are in two sides of the critical loop $\mathcal{C}$ (see Figure 5.2).
@@
Case 2: If $f$ is Type IIA, then $E_{0}\subsetneq\mathcal{C}\cap f(\mathcal{C})\subsetneq f(\mathcal{C})$. Similar as in Case 1, $f^{-1}(\mathcal{C})$ is a figure eight curve, and contains a shortest anchored simple closed curve, denoted by $\mathcal{C}_{1}$. Since $\mathcal{C}\cap f(\mathcal{C})$ contains more than $E_{0}$, $\mathcal{C}$ and $\mathcal{C}_{1}$ are siblings (see Figure 5.3). Similarly, $f^{-1}(\mathcal{C}_{1})$ contains one shortest anchored simple closed curve, denoted by $\mathcal{C}_{2}$. Note that $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$ are in two sides of the critical loop $\mathcal{C}$. Moreover $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$ are not siblings, and neither are $\mathcal{C}$ and $\mathcal{C}_{2}$. Then the same pull back argument as in Case 1 gives a sequence of shortest anchored simple closed curves $\mathcal{C}_{n}$, where $f(\mathcal{C}_{n})=\mathcal{C}_{n-1}$, and $\mathcal{C}_{n}$ has no sibling for all $n\geq 2$.
Case 3: If $f$ is Type IIB, then $\mathcal{C}\cap f(\mathcal{C})=f(\mathcal{C})$. Note that in this case, $\mathcal{C}$ contains both critical values. Thus, $f^{-1}(\mathcal{C})$ is a union of four simple paths connecting the critical points. Note that $f$ maps each one of the four components of $\hat{\mathbb{C}}-f^{-1}(\mathcal{C})$ homeomorphically to a component of $\hat{\mathbb{C}}-\mathcal{C}$. It is easy to verify that $f^{-1}(\mathcal{C})$ contains three shortest anchored simple closed curves, one of them is the critical loop. Denote them by $\mathcal{C}_{1},\mathcal{C}_{-1}$ and $\mathcal{C}$. Therefore, $\mathcal{C}$ has two siblings (see Figure 5.4).
Let $U_{1}$ and $U_{-1}$ be the regions bounded by $\mathcal{C}_{1}$ and $\mathcal{C}_{-1}$ that do not contain the critical loop $\mathcal{C}$, and let $\mathcal{V}_{1}$ and $\mathcal{V}_{-1}$ be the regions bounded by $\mathcal{C}$ that contains $\mathcal{C}_{1}$ and $\mathcal{C}_{-1}$ respectively. Then $f:U_{1}\longrightarrow V_{-1}$ and $f:U_{-1}\longrightarrow V_{1}$ are homeomorphisms. Thus, by inductively pulling back, we get a sequence of shortest anchored simple closed curves $\mathcal{C}_{n},n\in\mathbb{Z}$, where $f(\mathcal{C}_{n})=\mathcal{C}_{-\operatorname{sgn}(n)(|n|-1)}$ for $|n|\geq 1$. Here $\operatorname{sgn}(n)$ represents the sign of $n$, and $\mathcal{C}_{0}=\mathcal{C}$. Moreover, $\mathcal{C}_{n}$ has two siblings $\mathcal{C}_{n-1}$ and $\mathcal{C}_{n+1}$, and in addition $\mathcal{C}_{1}$ and $\mathcal{C}_{-1}$ are siblings. There are no other pairs of siblings, which gives the count in the proposition. ∎
@@ 6. Quasiconformal non-equivalence between quadratic gasket Julia set and limit set
A path $\gamma=[x_{0},...,x_{k}]$ in $\mathcal{G}$ is called a local geodesic if $x_{i},x_{j}$ are not adjacent in $\mathcal{G}$ if $|i-j|\geq 2$. Note that in particular, if $x,y$ are adjacent in $\mathcal{G}$, then there are no other local geodesics other than the edge $[x,y]$. We also remark that given any path $\gamma$, one can construct a local geodesic by replacing $x_{i},x_{i+1},...,x_{j}]$ with $x_{i},x_{j}]$ if $x_{i},x_{j}$ are adjacent in $\mathcal{G}$.
@@
  • $\gamma$ connects $a_{l}$ to $b_{i}$ for some i=1,,2l2i=1,...,2l-2;
@@
We claim that $\operatorname{Int}\gamma\cap f(\mathcal{C}_{0})=\emptyset$. Suppose not. Then we there exists a strictly shorter arc $\delta$ connecting $a_{l}$ to $v\in f(\mathcal{C}_{0})$. Note that $v\notin\{a,b\}$. Since $f$ maps the gap $R_{0}$ homeomorphically to its image, we can find a lift $\delta^{\prime}$ that connects $b_{l}$ and $a_{j}$ for some j=1,,2l2j=1,...,2l-2. Then $\tau(\delta^{\prime})$ is a local geodesic that connects $a_{l}$ to $b_{j}$ with $\operatorname{Int}\tau(\delta^{\prime})\subseteq\operatorname{Int}R_{0}$ and no interior vertex of $\tau(\delta^{\prime})$ is adjacent to $a_{0}$. This is a contradiction to the minimality of $\gamma$.
@@
separates the critical values $f(a_{l})$ and $b_{0}=f(a_{0})$ (see Figure 6.1). Since the arc $a_{0},a_{1}]\cup...\cup[a_{l-1},a_{l}]$ has a lift connecting $b_{l}$ and $b_{0}$, we conclude that the arc $\alpha$ has a lift connecting $b_{l}$ and $a_{1}$. ∎
@@
Suppose that $f(\beta)\nsubseteq\overline{R_{0}}$. Since $\beta$ connects $b_{l}$ to $b_{0}$ in $R_{0}$, and $f$ sends locally the region bounded between $b_{1},b_{0}]$ and $b_{0},a_{2l-2}]$ in $R_{0}$ to the region bounded by $a_{1},a_{0}]$ and $a_{0},f(a_{2l-2})]$ in $R_{0}$ (see Figure 6.1), we can decompose $f(\beta)$ as $f(\beta)=\gamma_{1}\cup\gamma_{2}\cup\gamma_{3}$ where $\gamma_{1}$ connects $a_{l}$ to $b_{i}$ for some $i$, $\gamma_{2}$ connects $b_{i}$ to $b_{j}$ for some $j$ and $\gamma_{3}$ connects $b_{j}$ to $a_{0}$. Moreover, $\operatorname{Int}\gamma_{1},\operatorname{Int}\gamma_{3}\subseteq\operatorname{Int}R_{0}$.
@@
[…]
@@
[…]
@@
By Proposition 5.3, any simple closed curve in $\mathcal{G}$ has length $\geq 2l$. Thus, for any two vertices $v,w$, there exists at most one path connecting $v,w$ with length <l<l. Moreover, if there is a path connecting $v,w$ with length $l$, then all the other paths have length $\geq l$. Since $b_{i},b_{j},a_{0}$ all lie on a shortest loop $\mathcal{C}_{1}$, we have that the lengths $l(\gamma_{2}),l(\gamma_{3})\geq l$. So